### v1

# Extended VC-dimension, and Radon and Tverberg type theorems for unions of convex sets

Abstract
We define and study an extension of the notion of the VC-dimension of a hypergraph and apply it to establish a Tverberg type theorem for unions of convex sets. We also prove a new Radon type theorem for unions of convex sets, vastly improving the estimates in an earlier result of Bárány and Kalai.
1 Introduction
Radon’s Theorem and Tverberg’s Theorem:
Radon’s theorem states that any set of $d+2$ points in $\mathbb{R}^{d}$ can be partitioned into two subsets whose convex hulls intersect. Formally, given a set $P=\{x_{1},x_{2},\ldots,x_{d+2}\}\subseteq\mathbb{R}^{d}$, there exists a partition of $P$ into two disjoint subsets $P_{1}$ and $P_{2}$ ($P=P_{1}\cup P_{2}$) such that $\text{CH}(P_{1})\cap\text{CH}(P_{2})\neq\emptyset$, where here and in what follows $CH(X)$ denotes the convex hull of a set of points $X$. The bound $d+2$ is tight as one can easily see by taking any set of $d+1$ affine-independent points in $\mathbb{R}^{d}$, that is, the vertices of the $d$-dimensional simplex.
This fundamental theorem is a cornerstone in discrete geometry, providing insights into the structure of point sets and their convex combinations. It implies various fundamental theorems in geometry including the classical Helly’s Theorem, the center point theorem, and more. Its implications extend to machine learning, statistical learning and computational geometry, influencing algorithms for geometric separation, point location, and convex hull computations. An example of such implication is the existence of small so-called ${\varepsilon}$-nets or ${\varepsilon}$-approximations for range-spaces defined by semi-algebraic sets which are a core notion in those areas. The combinatorial notion of VC-dimension is, in fact, the analog of Radon’s bound for abstract set-systems.
An equivalent formulation of Radon’s theorem states that for any set $P$ of $d+2$ points in $\mathbb{R}^{d}$, there exists a partition of the set into two subsets $P_{1}$ and $P_{2}$ such that any convex set containing $P_{1}$ must intersect any convex set containing $P_{2}$. This perspective highlights the strong separation properties that are impossible for sets of $d+2$ points in $\mathbb{R}^{d}$.
Moreover, Radon’s theorem serves as a basis for numerous generalizations and related results, such as the following classical and beautiful theorem of Tverberg [20], which further enriches our understanding of geometric configurations.
Theorem 1.1.
[Tverberg’s Theorem[20]] Let $r\geq 2$ be a fixed integer and $d\geq 1$. Then for any set $P$ of $(r-1)(d+1)+1$ points in $\mathbb{R}^{d}$ there exists a partition of $P$ into $r$ pairwise disjoint sets $P=\bigcup_{i=1}^{r}P_{i}$ such that $\bigcap_{i=1}^{r}\text{CH}(P_{i})\neq\emptyset$.
Note that Radon’s theorem is the special case of Tverberg’s theorem with $r=2$.
Tverberg’s Theorem has far-reaching implications in discrete and computational geometry and beyond. Combined with the colorful Carathéodory’s theorem of Bárány [5] it implies, for example, the so-called first selection Lemma (see, e.g., [15]).
An equivalent formulation of Tverberg’s theorem states that for any set $P$ of $(r-1)(d+1)+1$ points in $\mathbb{R}^{d}$, there exists a partition of the set into $r$ pairwise disjoint sets $P=\bigcup_{i=1}^{r}P_{i}$ such that for any family of convex sets $C_{1},\ldots,C_{r}$ with $P_{i}\subset C_{i}$ for every $i\in[r]$ we have that $\bigcap_{i=1}^{r}C_{i}\neq\emptyset$.
In this paper we study an old problem posed already in the 70’s by Gil Kalai [14] that asks for a generalization of Radon’s Theorem (and Tverberg’s Theorem) to sets which are not necessarily convex nor even connected, but are the union of a bounded number of convex sets.
Definition 1.2.
Let $s\geq 1$ be an integer. A set $C$ in $\mathbb{R}^{d}$ is said to be $s$-convex if it is the union of $s$ convex sets.
Problem 1.3.
[7] What is the least integer $f=f(d,s,t)$ such that for any set $P$ of $f$ points in $\mathbb{R}^{d}$ there is a partition $P=A\cup B$ such that any $s$-convex set containing $A$ must intersect any $t$-convex set containing $B$.
A more general problem is the following:
Problem 1.4.
What is the least integer $f=f_{r}(d,s_{1},\ldots,s_{r})$ such that for any set $P$ of $f$ points in $\mathbb{R}^{d}$ there is a partition into $r$ pairwise disjoint sets $P=\bigcup_{i=1}^{r}P_{i}$ such that for any family of sets $C_{1},\ldots,C_{r}$ with $P_{i}\subset C_{i}$ where $C_{i}$ is an $s_{i}$-convex set for every $i\in[r]$ we have that $\bigcap_{i=1}^{r}C_{i}\neq\emptyset$.
Notice that Radon’s theorem is equivalent to $f(d,1,1)=d+2$ and more generally Tverberg’s theorem is equivalent to $f_{r}(d,1,\ldots,1)=(r-1)(d+1)+1$.
Bárány and Kalai showed that $f(d,s,t)$ is always finite. However, their upper bound is a huge function of $d$, $s$ and $t$, due to their proof technique involving hypergraph Ramsey theory.
1.1 The main results
Our first main result is a fairly simple proof of the following near-optimal upper bound:
Theorem 1.5.
$f(d,s,t)=O(dst\log(st))$
Our second main result is an extension of the theorem above to the following version of Tverberg’s theorem for unions of convex sets and any $r\geq 2$:
Theorem 1.6.
$f_{r}(d,s_{1},\ldots,s_{r})=O(dr^{2}\cdot\log r\cdot\Pi_{i=1}^{r}s_{i}\cdot{\ln\Pi_{i=1}^{r}s_{i}})$
A notable feature of our proofs is that they are mainly combinatorial and therefore work for abstract separable convexity spaces with bounded Radon numbers:
Definition 1.7.
An abstract convexity space is a pair $(X,\mathcal{C})$ where $X$ is a set and $\mathcal{C}\subseteq 2^{X}$ is a family of subsets of $X$ called convex sets, satisfying the following axioms:
  1. 1.
    $X\in\mathcal{C}$ and $\emptyset\in\mathcal{C}$.
  2. 2.
    $\mathcal{C}$ is closed under intersections, i.e., if $\mathcal{D}\subseteq\mathcal{C}$, then $\bigcap\mathcal{D}\in\mathcal{C}$.
This generalizes the notion of convex sets $\mathbb{R}^{d}$. Abstract convexity spaces arise naturally in combinatorics, geometry, and lattice theory, and provide a unifying framework to study convexity. For more on abstract convexity spaces, see, e.g., [13, 16, 21].
For a subset $P\subset X$ in an abstract convexity space $(X,\mathcal{C})$ define its convex hull $CH(P)$ to be $CH(P)=\bigcap_{P\subset S,S\in\mathcal{C}}S$.
For a set $P\subset X$ we say that $P=P_{1}\cup P_{2}$ is a Radon partition if $CH(P_{1})\cap CH(P_{2})\neq\emptyset$. We say that an $r$ partition $P=\bigcup_{i=1}^{r}P_{i}$ is an $r$-Tverberg partition of $P$ if $\bigcap_{i=1}^{r}CH(P_{i})\neq\emptyset$. The Radon number $r(X,\mathcal{C})$ of a space $(X,\mathcal{C})$ is the minimum integer $n$ such that any subset $P\subset X$ with $\lvert{P}\rvert\geq n$ admits a Radon partition. If $n$ is unbounded then the Radon number is $\infty$. Similarly, the $r$’th Tverberg number $T_{r}(X,\mathcal{C})$ is the minimum integer $n$ such that any subset $P\subset X$ with $\lvert{P}\rvert\geq n$ admits an $r$-Tverberg partition.
A convex set $H\in\mathcal{C}$ in an abstract convexity space $(X,\mathcal{C})$ is called a halfspace if its complement is also a convex set i.e., $X\setminus H\in\mathcal{C}$.
We say that an abstract convexity space $(X,\mathcal{C})$ is separable if for any two disjoint sets there exists a separating halfspace. Formally, for any $C_{1}\in\mathcal{C}$ and $C_{2}\in\mathcal{C}$ if $C_{1}\cap C_{2}=\emptyset$ then there exists a halfspace $H$ such that $C_{1}\subset H,C_{2}\subset X\setminus H$.
Our proof technique of Theorem 1.5 and Theorem 1.6 works for abstract convexity spaces which are separable and have a bounded Radon number. Let $(X,\mathcal{C})$ be a separable abstract convexity space with Radon number $d=r(X,\mathcal{C})$. Call a subset $A\subset X$ $s$-convex if it is the union of $s$ convex sets in $\mathcal{C}$. Let $F=F(d,s,t)$ be the minimum integer so that any set $P$ of cardinality at least $F$ in a separable abstract convexity space $(X,\mathcal{C})$ with Radon number $d$ admits a partition $P=P_{1}\cup P_{2}$ such that any $s$-convex set containing $P_{1}$ must intersect any $t$-convex set containing $P_{2}$. We have:
Theorem 1.8.
$F(d,s,t)=O(dst\log{st})$.
More generally let $F=F_{r}(d,s_{1},\ldots,s_{r})$ be the minimum integer so that any set $P$ of cardinality at least $F$ in a separable abstract convexity space $(X,\mathcal{C})$ with Radon number $d$ admits a partition $P=\bigcup_{i=1}^{r}P_{i}$ into $r$ pairwise disjoint sets $P=\bigcup_{i=1}^{r}P_{i}$ such that for any family of sets $C_{1},\ldots,C_{r}$ with $P_{i}\subset C_{i}$ where $C_{i}$ is an $s_{i}$-convex set for every $i\in[r]$ we have that $\bigcap_{i=1}^{r}C_{i}\neq\emptyset$.
Theorem 1.9.
$F_{r}(d,s_{1},\ldots,s_{r})=O(dr^{2}\cdot\log r\cdot\Pi_{i=1}^{r}s_{i}\cdot{\ln\Pi_{i=1}^{r}s_{i}})$.
We describe here the proofs of Theorem 1.5 and Theorem 1.6. Essentially the same proofs establish the corresponding results, Theorem 1.8 and Theorem 1.9, for abstract separable convexity spaces.
1.2 Tools and preliminaries
Our proof of Theorem 1.5 is based on some simple properties of the VC-dimension and the shatter function of the relevant hypergraphs. In order to prove Theorem 1.6 we define an extended version of the VC-dimension and show how to bound it in terms of the usual dimension.
The Vapnik-Chervonenkis dimension of a hypergraph is a measure of its complexity, which plays a central role in statistical learning, computational geometry, and other areas of computer science and combinatorics (see, e.g., [3, 2, 4, 6, 17]). Many graphs and hypergraphs that arise in geometry have bounded VC-dimension.
Definition 1.10 (VC-dimension).
The Vapnik-Chervonenkis dimension $VC(H)$ of a hypergraph $H=(V,E)$ is the largest integer $d$ such that there exists a subset $S\subseteq V$ (not necessarily in $E$) with $|S|=d$ that is shattered by $E$. A subset $S$ is said to be shattered by $E$ if, for every subset $T\subseteq S$, there exists a hyperedge $e\in E$ such that $e\cap S=T$.
Note that by the fact that every two disjoint convex sets in $\mathbb{R}^{d}$ are separable by a halfspace, Radon’s Theorem implies that no set of $d+2$ points in $\mathbb{R}^{d}$ is shattered with halfspaces. Namely, if ${\cal H}_{d}$ is the family of all halfspaces in $\mathbb{R}^{d}$ then the VC-dimension of the hypergraph $(\mathbb{R}^{d},{\cal H}_{d})$ is at most $d+1$. This simple observation holds for every separable abstract convexity space $(X,\mathcal{C})$ with Radon number $d$. In that case the VC-dimension of the hypergraph $(X,\cal H)$ where $\cal H$ is the family of all halfspaces in $\mathcal{C}$ is at most $d-1$.
Definition 1.11.
The primal shatter function of a hypergraph $H=(V,E)$ is the following function $\pi_{H}:\mathbb{N}\rightarrow\mathbb{N}$:
$\pi_{H}(m)=\max_{S\subseteq V,|S|=m}|\{S\cap e:e\in E\}|.$
The value $\pi_{H}(m)$ represents the maximum number of distinct subsets of a set $S$ of cardinality $m$ that can be realized as intersections with hyperedges in $E$.
The following lemma known as the Sauer-Perles-Shelah lemma provides an upper bound on the shatter function for hypergraphs with bounded VC-dimension (See, e.g., [15]):
Lemma 1.12 (Sauer-Perles-Shelah).
Let $H=(V,E)$ be a hypergraph with VC dimension $d$. Then
$\pi_{H}(m)\leq\sum_{i=0}^{d}\binom{m}{i}.$
In particular, if $m>d$, then $\pi_{H}(m)\leq(\frac{em}{d})^{d}$.
In order to tackle Problem 1.4 we need to develop an analogous notion of a shattered set in hypergraphs with bounded VC-dimension to partitions with more than $2$ parts. The relevant definition follows.
Definition 1.13.
Let $H=(V,E)$ be a fixed hypergraph. A subset $S\subset V$ is said to be $r$-shattered by $E$ if for any partition of $S$ into $r$ pairwise disjoint sets $S=\bigcup_{i=1}^{r}S_{i}$ there exists hyperedges $e_{1},\ldots,e_{r}\in E$ such that $S_{i}\subset e_{i}$ for all $i\in[r]$ and $S\cap\bigcap_{i=1}^{r}e_{i}=\emptyset$.
The following combinatorial lemma provides an extension of the Sauer-Perles-Shelah Lemma and might be of independent interest:
Lemma 1.14.
Let $d$ be a fixed integer and let $H=(V,E)$ be a hypergraph with VC-dimension $d$. Let $r\geq 2$ be an integer. Suppose that
$\big{(}\sum_{i=0}^{d}{f\choose i}\big{)}^{r}<(\frac{r}{r-1})^{f}$ (1)
Then any subset $S$ of $f$ vertices cannot be $r$-shattered. Namely, there exists a partition $S=\bigcup_{i=1}^{r}S_{i}$ so that whenever we have $r$ hyperedges $e_{1},e_{2},\ldots,e_{r}\in E$ such that $S_{i}\subset e_{i}$ for all $i\in[r]$ it must hold that $S\cap\bigcap_{i=1}^{r}e_{i}\neq\emptyset$.
Note that the inequality (1) holds for $f=Cdr^{2}\log r$ for some absolute constant $C$.
The proof of the lemma is described in Section 3.
1.3 Structure
The rest of this short paper is organized as follows. In Section 2 we describe the proof of Theorem 1.5. The proof of Theorem 1.6 is given in Section 3. After a discussion of lower and upper bounds for the relevant functions for various ranges of the parameters in Sections 4 and 5 we suggest several open problems in the final Section 6.
2 Proof of Theorem 1.5
Before proceeding to the proof of Theorem 1.5 we need the following two easy Lemmas.
Lemma 2.1.
Let $C_{1}$ be an $s$-convex set in $\mathbb{R}^{d}$ and $C_{2}$ a $t$-convex set. Assume that $C_{1}\cap C_{2}=\emptyset$. Then there exist $s$ convex polytopes $K_{1},\ldots,K_{s}$ with a total of at most $st$-facets whose union covers $C_{1}$ so that the complement of the union covers $C_{2}$. Namely, $C_{1}\subset\bigcup_{i=1}^{s}K_{i}$ and $C_{2}\subset\overline{\bigcup_{i=1}^{s}K_{i}}$.
Proof.
Since $C_{1}$ is $s$-convex it can be written as $C_{1}=\bigcup_{i=1}^{s}X_{i}$ for some convex sets $X_{1},\ldots,X_{s}$. Similarly $C_{2}=\bigcup_{j=1}^{t}Y_{i}$ for $t$ convex sets $Y_{1},\ldots,Y_{t}$. Since $C_{1}\cap C_{2}=\emptyset$, for every $i\in[s],j\in[t]$ we have that $X_{i}\cap Y_{j}=\emptyset$ and therefore there exists a hyperplane $h_{i,j}$ strictly separating $X_{i}$ and $Y_{j}$. Assume without loss of generality that the positive open halfspace $h_{i,j}^{+}$ bounded by $h_{i,j}$ contains $X_{i}$ and the negative open halfspace $h_{i,j}^{-}$ contains $Y_{j}$. For every $i\in[s]$ let $K_{i}$ be the convex polytope which is the intersection $\bigcap_{j=1}^{t}h_{i,j}^{+}$. Note that $K_{i}$ is a convex polytope with at most $t$ facets containing $X_{i}$ and its complement $\overline{K_{i}}$ contains $C_{2}$. So the union of the polytopes $\bigcup_{i=1}^{s}K_{i}$ contains $C_{1}$ and its complement $\overline{\bigcup_{i=1}^{s}K_{i}}$ contains $C_{2}$. Moreover, the total number of facets of these polytopes is at most $st$. This completes the proof of the lemma. ∎
Lemma 2.2.
Let $l$ be an integer and let $H=(P,E)$ be a hypergraph where $P$ is a set of points in $\mathbb{R}^{d}$ and $S\in E$ is a hyperedge if and only if there exists a set $K_{1},\ldots,K_{i}$ of convex polytopes with a total of at most $l$ facets such that $S=P\cap(\bigcup_{j=1}^{i}K_{j})$. Namely, $S$ can be cutoff from $P$ by intersecting it with a set consisting of a union of convex polytopes with a total of $l$ facets. Then the VC-dimension of $H$ is bounded by $O(dl\log l)$.
Proof.
The proof is rather standard and follows from e.g., [15](Proposition 10.3.3). ∎
Proof of Theorem 1.5: Put $l=st$. Let $H=(\mathbb{R}^{d},E)$ be the hypergraph as in Lemma 2.2. Let $n=O(dl\log l)=O(dst\log(st))$ be its VC-dimension. We claim that $f(d,s,t)\leq n+1$. Indeed, Let $P$ be a set of $n+1$ points in $\mathbb{R}^{d}$. Since $P$ cannot be shattered by the hyperedges in $H$ there exists a non-trivial subset $A\subset P$ such that no hyperedge $S\in E$ has the property that $S\cap P=A$. In other words there does not exists a set $K$ which is the union of convex polytopes with a total of at most $l$ facets such that $K\cap P=A$. We claim that the partition $P=A\cup(P\setminus A)$ has the property that every $s$-convex set containing $A$ must intersect any $t$-convex set containing $P\setminus A$. Indeed, assume to the contrary that there exists an $s$-convex set $C_{1}$ containg $A$ and a $t$-convex set $C_{2}$ containing $P\setminus A$ such that $C_{1}\cap C_{2}=\emptyset$. Then by Lemma 2.1 there exists a set $K$ which is the union of convex polytopes with a total of at most $l=st$ facets containing $C_{1}$ with complement $\overline{K}$ containing $C_{2}$. In particular, $K$ contains $A$ and $\overline{K}$ contains $P\setminus A$ so $K\cap P=A$, a contradiction. This completes the proof. $\Box$
3 Generalized Tverberg’s Theorem
In this section we tackle Problem 1.4. In what follows we provide a bound on $f_{r}(d,s_{1},\ldots,s_{r})$. For the sake of simplicity of computations, we assume that $s_{1}=s_{2}=\cdots=s_{r}=s$ and abuse the notation writing $f_{r}(d,s)$ for $f_{r}(d,s,s,\ldots,s)$. Our proof technique can be easily modified to make the bound sensitive to any $r$ integer parameters $s_{1},\ldots,s_{r}$ in the more general setting.
Our argument is based on the notion defined in 1.13, which is an extension of the notion of VC-dimension to partitions with more than $2$ sets.
We need Lemma 1.14. We proceed with its proof.
Proof.
Suppose the statement of the lemma is false and there is a set $S$ that violates the condition. For every (ordered) $r$-tuple of hyperedges $e_{1},e_{2},\ldots,e_{r}$ with no common intersection in $S$, every point of $S$ belongs to at most $r-1$ of these hyperedges. Therefore, these fixed $r$ hyperedges can be used to provide at most $(r-1)^{f}$ partitions into $r$ sets $S_{1},\ldots,S_{r}$. Indeed every point among the $f$ points of $S$ has at most $r-1$ options to decide to which part of the partition it belongs- if, for example, a point lies in all hyperedges $e_{1}$ and $e_{2}$ up to $e_{r-1}$ it can be in the partition in either $S_{1}$ or $S_{2}$ up to $S_{r-1}$ but not in $S_{r}$, and similarly for each other case. By the Sauer-Perles-Shelah Lemma 1.12, the number of ordered $r$-tuples of intersections of hyperedges with $S$ is at most
$\left(\sum_{i=0}^{d}{f\choose i}\right)^{r}$
Since we have to cover all $r^{f}$ ordered partitions of $S$ into $r$ parts we get
$\left(\sum_{i=0}^{d}{f\choose i}\right)^{r}\cdot(r-1)^{f}\geq r^{f}$
This contradicts the assumption (1) and completes the proof. ∎
We also need the following simple geometric result.
Lemma 3.1.
Let $C_{1},C_{2},\ldots,C_{r}$ be $r$ sets in $\mathbb{R}^{d}$ where each set is an $s$-convex set. Assume that $\bigcap_{i=1}^{r}C_{i}=\emptyset$. Then there exist $r$ sets $K_{1},\ldots,K_{r}$ where each $K_{i}$ is the union of $s$ convex polytopes with a total of at most $s^{r}$-facets such that $C_{i}\subset K_{i}$ for all $i\in[r]$ and $\bigcap_{i=1}^{r}K_{i}=\emptyset$.
Proof.
Since for any $i\in[r]$ $C_{i}$ is $s$-convex it can be written as $C_{i}=\bigcup_{j=1}^{s}X_{i,j}$ for some convex sets $X_{i,1},\ldots,X_{i,s}$. Since $\bigcap_{i=1}^{r}C_{i}=\emptyset$, then for every $i\in[r]$ we have that $C_{i}$ is disjoint from $B_{i}=\bigcap_{j\neq i}C_{j}$. Note that each such $B_{i}$ is the union of at most $s^{r-1}$ convex sets since it is an $(r-1)$-fold intersection of unions of $s$-convex sets. We construct the sets $K_{1},\ldots,K_{r}$ one by one. First, we replace $C_{1}$ by $K_{1}$ separating it from the union of at most $s^{r-1}$ convex sets $B_{1}=\bigcap_{j>1}C_{j}$. As before, this can be done with $K_{1}$ which is the union of $s$ convex polytopes with a total of at most $s^{r}$ facets. In particular $K_{1}$ is also $s$-convex. Also $C_{1}\subset K_{1}$. Moreover, $K_{1}\cap\bigcap_{j>1}C_{j}=\emptyset$. We then apply the same argument to $C_{2}$ as the intersection of $K_{1},C_{2},..,C_{r}$ is empty and all the sets are $s$-convex. So we can find a set $K_{2}$ which is $s$-convex and consists of the union of $s$ convex poytopes with a total of at most $s^{r}$ facets and such that $C_{2}\subset K_{2}$ and $K_{1}\cap K_{2}\cap\bigcap_{j>2}C_{j}=\emptyset$. Continuing in the same manner we conclude that each $C_{i}$ can be replaced with such a $K_{i}$ so that $C_{i}\subset K_{i}$ for all $i\in[r]$, each $K_{i}$ is the union of $s$ convex polytopes with a total of $s^{r}$ factes and $\bigcap_{i=1}^{r}K_{i}=\emptyset$. This completes the proof of the lemma. ∎
We are now ready to prove the following theorem extending Theorem 1.5 to partitions with $r>2$ parts, and generalizing Tverberg’s theorem to $s$-convex sets:
Proof of Theorem 1.6 (for $s_{1}=s_{2}=\ldots=s_{r}=s$).
Put $l=s^{r}$. Let $H=(\mathbb{R}^{d},E)$ be the hypergraph as in Lemma 2.2. Let $d^{\prime}=O(dl\log l)$ be its VC-dimension. Let $n$ be the maximum size of an $r$-shattered set. Note that by Lemma 1.14 $n=O(d^{\prime}r^{2}\log r)=O(dr^{2}{\log r}\cdot s^{r}\cdot{\log s^{r}})$
We claim that $f_{r}(d,s)\leq n+1$. Indeed, Let $P$ be a set of $n+1$ points in $\mathbb{R}^{d}$. Since $P$ cannot be $r$-shattered by the hyperedges in $H$ there exists a partition $P=\bigcup_{i=1}^{r}P_{i}$ such that whenever we have $r$ hyperedges $e_{1},\ldots,e_{r}\in E$ with $P_{i}\subset e_{i}$ for each $i\in[r]$ it follows that $P\cap\bigcap_{i=1}^{r}e_{i}\neq\emptyset$. In other words there does not exists a sets $K_{1},\ldots,K_{r}$ each of which is the union of convex polytopes with a total of at most $s^{r}$ facets such that $P_{i}\subset K_{i}$ for every $i\in[r]$ and $\bigcap_{i=1}^{r}K_{i}=\emptyset$. We claim that the partition $P=\bigcup_{i=1}^{r}P_{i}$ has the property that for every family of $r$ sets $C_{1},\ldots,C_{r}$ such that for each $i\in[r]$ $P_{i}\subset C_{i}$ and each $C_{i}$ is an $s$-convex set it must hold that $\bigcap_{i=1}^{r}C_{i}\neq\emptyset$. Indeed, assume to the contrary that there exist $C_{1},\ldots,C_{r}$ such that for each $i\in[r]$ $P_{i}\subset C_{i}$ and each $C_{i}$ is an $s$-convex set and such that $\bigcap_{i=1}^{r}C_{i}=\emptyset$. Then by Lemma 3.1 there exist sets $K_{1},\ldots,K_{r}$ and $C_{i}\subset K_{i},\forall i\in[r]$ and each $K_{i}$ is the union of convex polytopes with a total of at most $l=s^{r}$ facets and $\bigcap_{i=1}^{r}K_{i}=\emptyset$, a contradiction. This completes the proof. ∎
4 A lower bound for $f_{r}(d,s)$
An easy lower bound for the function $f_{r}(d,s)$ can be proved by taking $s$ translated copies of an extremal example for the classical theorem of Tverberg with pairwise disjoint convex hulls. This gives the following.
Proposition 4.1.
$f_{r}(d,s)>s(d+1)(r-1)$
This is, of course, tight for $s=1$, as Tverberg’s Theorem is tight. In this section we show that, somewhat surprsingly, for $s\geq 3$ the lower bound becomes quadratic in $r$. As we show in the next section, this is the case even in dimension $d=1$, where the lower bound is tight up to a constant factor for all $r$ and $s\geq 3$. For convenience we describe the proof for even $r$, a similar bound for odd $r$ follows from the one for $r-1$.
Theorem 4.2.
For every $d\geq 1,s\geq 3$ and even $r\geq 2$,
$f_{r}(d,s)>\frac{1}{4}(\lfloor\frac{d}{2}\rfloor+1)\lfloor\frac{s-1}{2}\rfloor r^{2}$
Proof.
Put $m=(\lfloor\frac{d}{2}\rfloor+1)r/2$, $p=\lfloor\frac{s-1}{2}\rfloor r/2$ and $n=mp$. Let $P$ be a set of $n$ points on the moment curve in $\mathbb{R}^{d}$ consisting of the $n$ points $z_{i}=(t_{i},t_{i}^{2},\ldots t_{i}^{d})$, where $10<t_{1}<t_{2}<\ldots<t_{n}<$ 1. Let $I_{1},I_{2},\ldots,I_{p}$ be $p$ open intervals that cover $(0,1)$ and appear on it in this order, where the left endpoint of each interval $I_{j+1}$ is just slightly smaller than the right endpoint of $I_{j}$ for each $j$. The intervals are chosen so that each $t_{i}$ belongs to exactly one of them, and each interval contains exactly $m$ points $t_{j}$.
In order to complete the proof we show that for any coloring of the points of $P$ by $r$ colors there are $s$-convex sets $C_{i}$, with $C_{i}$ containing all points of color $i$, so that the intersection of all the sets $C_{i}$ is empty. Each set $C_{i}$ will be defined as the union of at most $s$ convex sets, where each of the convex sets will be the convex hull of points of color $i$ with first coordinate lying in a union of some consecutive intervals $I_{q}$. The crucial property of the definition of these convex sets is that for each interval $I_{q}$ there will be an index $i$ so that one of the convex sets in the definition of $C_{i}$ will be the set of all points $z_{j}=(t_{j},t_{j}^{2},\ldots,t_{j}^{d})$ of color $i$ for which $t_{j}$ lies in this single interval $I_{q}$. Moreover, this will be done in a way that ensures that the number of such points is always at most $\lfloor d/2\rfloor$.
Note that if such a choice indeed exists, then the intersection of the corresponding $r$ sets $C_{i}$ will be empty as needed. Indeed, otherwise the intersection is some point $z\in\mathbb{R}^{d}$ (not necessarily on the moment curve). Let $t$ denote the first coordinate of this point $z$. Suppose first that $t$ belongs only to one interval $I_{q}$, and let $i$ be the color chosen for $I_{q}$ so that there are at most $\lfloor d/2\rfloor$ points of color $i$ with their first coordinate in $I_{q}$. Then $z$ has to lie in the convex hull of these points (as any other convex set among the ones defining $C_{i}$ either has all its points with first coordinate smaller than $t$ or all its points with first coordinate larger than $t$). Since the moment curve is $\lfloor d/2\rfloor$-neighborly, this convex hull is disjoint from the convex hull of all the other points of $P$, implying that $z$ cannot lie in any other set $C_{g}$ besides $C_{i}$. If $t$ belongs to two intervals, say $I_{q-1}$ and $I_{q}$ and $i$ is defined as before, then $z$ cannot lie in $C_{i}$, since in this case any convex set among those defining $C_{i}$ either has all its points with first coordinate smaller than $t$ or all its points with first coordinate larger than $t$.
It thus only remains to show that we can choose for every interval $I_{q}$ a color $i$ with the required properties, making sure that no fixed color is chosen more than $\lfloor(s-1)/2\rfloor$ times (as this way each set $C_{i}$ will be the union of at most $s$ convex sets). To do so we go over the intervals $I_{q}$ one by one, in an arbitrary order (for example, from left to right). When dealing with the interval $I_{q}$ after handling several previous ones, we first note that since there are exactly $m=(\lfloor d/2\rfloor+1)r/2$ points with first coordinate in $I_{q}$, there are at least $r/2$ colors that appears at most $\lfloor d/2\rfloor$ times in $I_{q}$. As so far we have selected the colors $i$ for at most $p-1=\lfloor(s-1)/2\rfloor r/2-1$ intervals, there are less than $r/2$ colors $i$ that have already been chosen $\lfloor(s-1)/2\rfloor$ times. It follows that there is at least one color $i$ that can be chosen for the interval $I_{q}$, implying that the process terminates successfuly. This completes the proof of the theorem. ∎
5 Improved bounds in special cases
In this section we establish improved upper and lower bounds for the functions $f$ and $f_{r}$ in several special cases.
Theorem 5.1.
$f(2,s,1)=2s+2$
and for every $d\geq 4$
$f(d,s,1)=\Theta(ds\log s)$
Proof.
Let $l$ denote the VC-dimension of a hypergraph defined by points in $\mathbb{R}^{d}$ with respect to convex polytopes with at most $s$ facets. We first show that $f(d,s,1)=l+1$. Combined with the known bounds on the VC-dimension of such hypergraphs we get the claimed bounds. If $l$ is the VC-dimension of such a hypergraph then there exists a shattered set $P$ of size $l$. Then for any parition $P=P_{1}\cup P_{2}$ there is a polytope $K$ with at most $s$ facets such that $P_{2}\subset K$ and $K\cap P_{1}=\emptyset$. In particular $P_{1}$ is contained in the union of the at most $s$ complement halfspaces supporting the facets of $K$, which is an $s$-convex set (consisting of the at most $s$ complement halfspaces), and $P_{2}$ lies in its complement. Since this holds for any partition it follows that $f(d,s,1)\geq l+1$. In order to show equality we need to show that every $l+1$ point set in $\mathbb{R}^{d}$ admits a partition that cannot be realized by intersections with disjoint convex and $s$-convex sets. This follows by the same argument from the fact that any set $P$ of $l+1$ points cannot be shattered in the corresponding hypergraph and therefore there is at least one partition $P=P_{1}\cup P_{2}$ so that no polytope with at most $s$ facets can contain $P_{2}$ while being disjoint from $P_{1}$. Hence any $s$-convex set $C$ containing $P_{1}$ must intersect the convex hull $CH(P_{2})$ for otherwise by separation arguments as above there would be a polytope which is the intersection of $s$ half spaces containing $P_{2}$ and disjoint from $P_{1}$, a contradiction. It is a simple exercise to see that in the plane ($d=2$), the VC-dimension is $l=2s+1$ and thus $f(2,s,1)=l+1=2s+2$. The statement for any fixed $d\geq 4$ follows from the results in [10]. ∎
Theorem 5.2.
$f(2,s,s)\geq 4s$
Proof.
Let $P$ be a set of $14s-$ 1 points placed along a cycle $U$ in the plane. We have to show that for any partition of $P$ into two disjoint sets $A$ and $B$ there are disjoint $s$-convex sets $C_{1}$ and $C_{2}$ so that $A\subset C_{1}$ and $B\subset C_{2}$. Partition the set of points $P$ into disjoint subsets, where each subset is a maximal subset of $P$ consisting of consecutive points along $U$ that are all in $A$ or all in $B$. Let this partition be $A_{1},B_{1},A_{2},B_{2},\ldots,A_{t},B_{t}$, where each $A_{i}$ is a subset of $A$ and each $B_{i}$ is a subset of $B$, and the sets appear in this order along $U$. Clearly $t\leq 2s-1$, since each of the sets $A_{i},B_{j}$ is nonempty. If $t<2s-1$ define $A_{j}=B_{j}=\emptyset$ for all $t<j\leq 2s-1$. Let $C_{1}$ consist of the union of the following $s$ convex sets (some of which may be empty): the convex hull of $A_{1}\cup A_{2}\cup\ldots\cup A_{s}$, the convex hull of $A_{s+1}$, the convex hull of $A_{s+2}$, $\ldots$, the convex hull of $A_{2s-1}$. Then $C_{1}$ is $s$-convex and contains $A$. Similarly, let $C_{2}$ be the union of the following $s$ convex sets: the convex hull of $B_{s}\cup B_{s+1}\cup\ldots\cup B_{2s-1}$, the convex hull of $B_{1}$, the convex hull of $B_{2}$, $\ldots$, the convex hull of $B_{s-1}$. Then $C_{2}$ is clearly $s$-convex and contains $B$. It is easy to check that $C_{1}$ and $C_{2}$ are disjoint, completing the proof. ∎
Theorem 5.3.
$f(3,s,1)\leq 4s+1$
Proof.
We need to prove that for every set $P$ of $14s+$ 1 points there exists a partition $P=A\cup B$ such that any convex set containing $A$ must intersect any $s$-convex set containing $B$. Notice that by the above arguments it is enough to prove that the VC-dimension of the hypergraph $H=(P,E)$ where $E$ is the family of all intersections of $P$ with a convex polytope with at most $s$ facets is bounded by $4s$. This fact is proved in [11]. For completeness we include a proof based on an argument in [19]. Assume to the contrary there there exists a set $P\subset\mathbb{R}^{3}$ of size $14s+$ 1 that is shattered by $H$. In particular, for any partition $P=A\cup B$ if $A$ can be separated from $B$ by a convex polytope with at most $s$ facets, then there exists a set of $s$ half spaces whose union contains $B$ but none of the points in $A$.
Next, we need the following fact which was proved in [19]: There exists a $4$-coloring of the points of $P$ such that no halfspace that contains at least two points of $P$ is monochromatic.
Consider such a coloring. By the pigeonhole principle there is a monochromatic set $B\subset P$ of size at least $s+1$. We claim that $B$ cannot be separated from $A=P\setminus B$ with a union of only $s$ halfspaces. Indeed, if such $s$ halfspaces exist then one of them must contain at least $2$ points of $B$ and none of the points in $A$ so such a halfspace cuts off a monochromatic set of points, a contradiction. This completes the proof.
We next show that $g(3,s,s)$ is super-linear in $s$.
Theorem 5.4.
There exists a function $g(s)$ tending to infinity as $s$ tends to infinity so that
$f(3,s,s)\geq sg(s)$
Proof.
We make no attempt to optimize the function $g(s)$, and only show that it can be chosen as a function tending to infinity with $s$. Following the approach in [1], the proof applies the result of Furstenberg and Katznelson [12] known as the density Hales-Jewett Theorem. For an integer $k\geq 2$, put $[k]=\{1,2,\ldots,k\}$ and let $[k]^{d}$ denote the set of all vectors of length $d$ with coordinates in $[k]$. A combinatorial line is a subset $L\subset[k]^{d}$ so that there is a set of coordinates $I\subset[d]=\{1,2,\ldots,d\}$, $I\neq[d]$, and values $k_{i}\in[k]$ for all $i\in I$ for which $L$ is the following set of $k$ members of $[k]^{d}$:
$L=\{\ell_{1},\ell_{2},\ldots,\ell_{k}\}$
where
$\ell_{j}=\{(x_{1},x_{2},\ldots,x_{d})~{}:~{}x_{i}=k_{i}~{}\mbox{for all}~{}i\in I~{}\mbox{and}~{}x_{i}=j~{}\mbox{for all}~{}i\in[d]\setminus I\}.$
Thus a combinatorial line is a set of $k$ vectors all having some fixed values in the coordinates in $I$, where the $j$th vector has the value $j$ in all other coordinates. In this notation, the Furstenberg-Katznelson Theorem is the deep result that for any fixed integer $k$ and any fixed $\delta>0$ there exists an integer $d_{0}=d_{0}(k,\delta)$ so that for any $d\geq d_{0}$, any set $Y$ of at least $\delta k^{d}$ members of $[k]^{d}$ contains a combinatorial line.
Fix a small positive real $\delta$, a large integer $k$ and a huge integer $d=d_{0}(k,\delta)$ defined as above. View the points of $[k]^{d}$ as points in the $d$-dimensional real space $\mathbb{R}^{d}$ and call a (geometric) line in this space special if it contains all $k$ points of a combinatorial line.
By the claim in the proof of Theorem 1.3 in [1] the only points that belong to at least two special lines are the points of $[k]^{d}.$ Equivalently, if two such lines do not have a common point of $[k]^{d}$ then the full geometric lines are disjoint.
Now project all the configuration of the $n=k^{d}$ points above randomly to the $3$-dimensional space $\mathbb{R}^{3}$. This gives a set $P$ of $n$ points in $\mathbb{R}^{3}$. With probability $1$, the condition about the intersection of the projected special lines still holds: if two of them do not have a point of $P$ in common, then they are disjoint.
Fix a partition of $P$ into two disjoint sets $A$ and $B$. The following procedure partitions each of these sets into less than $\delta k^{d}+k^{d-1}$ pairwise disjoint subsets, which also have pairwise disjoint convex hulls.
Starting with the full set $A$, as long as it contains at least $\delta k^{d}$ points choose a combinatorial line in it. Define a subset consisting of the projected images of the points of this line, and remove all these points from $A$. Once the remaining size of $A$ is smaller than $\delta k^{d}$ take every single point as a subset. Handle $B$ in the same way. Defining $s$ as $s=\delta k^{d}+k^{d-1}$, this shows that $f(3,s,s)>n=k^{d}$. Since $\delta$ is arbitrarily small and $k$ is arbitrarily large, this shows that the ratio $g(s)=f(3,s,s)/s$ tends to infinity as $s$ tends to infinity, completing the proof. It is worth noting that the estimate for the growth of $g(x)$ can be improved using the same reasoning together with the results in [8], but this will still leave a large gap between the upper and lower bounds we know for $f(3,s,s)$. ∎
The following simple result shows that the lower bound for $f_{r}(d,s)$ proved in Section 4 is tight up to a constant factor in dimension $d=1$.
Theorem 5.5.
$f_{r}(1,s)\leq r(r-1)(s+1)+1$
Proof.
Put $n=r(r-1)(s+1)+1$ and let $10<p_{1}<p_{2}<\ldots<p_{n}<$ 1 be a set of $n$ points on the line. We have to show that there is a coloring of these points by $r$ colors $10,1,2,\ldots,r-$ 1, so that for any $r$ sets $C_{i}$, where each $C_{i}$ is a union of at most $s$ intervals that covers all points of color $i$, there is a point that lies in all sets $C_{i}$. Naturally, the coloring we choose colors the points periodically, that is, the color of $p_{i}$ is defined to be $i\bmod r$. Let $C_{i}$ be collections of intervals as above. Note that each $C_{i}$ must contain all points of $P$ besides at most $(s+1)(r-1)$. Indeed, since it contains all points colored $i$, the gap between any two consecutive intervals in it contains at most $r-1$ points, and the same holds for the gap between 00 and its leftmost point and between $1$ and its rightmost point. (We note that here we can slightly improve the bound since, for example, the gap between 00 and the leftmost point of $C_{i}$ can contain at most $i$ points $p_{i}$). It follows that if $n>r(r-1)(s+1)$ then there is a point (of $P$, although that’s not needed) that belongs to all sets $C_{i}$. This completes the proof. ∎
6 Concluding remarks and open problems
We established extensions of Radon’s Theorem and Tverberg’s Theorem for unions of convex sets. The main tools in the proofs are upper bounds for the shatter functions of range spaces with a bounded VC-dimension as well as an extension of these results. This extension, defined and studied here, is useful in the study of partitions with more than $2$ parts, which are the ones considered in the classical definition of the VC-dimension.
Our Theorem 1.9 for the special case $s=1$ provides an upper bound of $O(dr^{2}\log r)$ on the so-called Tverberg number in separable abstract convexity spaces in terms of its Radon number $d$. This improves the upper bound of $c(d)r^{2}\log^{2}r$ by Bukh [9] who proved it for the more general setting of (not necessarily separable) abstract convexity spaces. In [18] Pálvölgyi provided an upper bound of the form $O(d^{d^{d^{\log d}}}r)$ which is linear in $r$ but super exponential in $d$. It will be interesting to decide if one can get rid of the separability assumption,
While our upper and lower bounds for the functions $f(d,s,t)$ and $f_{r}(d,s_{1},s_{2},\ldots,s_{r})$ proved here are not very far from each other, the problem of determining them precisely remains open for most values of the parameters. It will be interesting to close the gap between the upper and lower bounds. The following specific questions are particularly intriguing.
Problem 6.1.
Is $f_{2}(s,s)$ linear in $s$
Problem 6.2.
Is $f_{r}(d,s)$ upper bounded by a polynomial in $r$,$d$ and $s$
Acknowledgements
We thank Gil Kalai for helpful discussions.