### v3

# arXiv:2506.17777v3[math.CO] 27 Feb 2026

Abstract
We prove a new Radon type theorem for unions of convex sets, settling an open problem posed by Kalai in the 1970s. We also define and study an extension of the notion of the VC-dimension of a hypergraph and apply it to establish an extension of our Radon type theorem to a Tverberg type theorem for unions of convex sets.
1 Introduction
Radon’s theorem states that any set of $d+2$ points in $\mathbb{R}^{d}$ can be partitioned into two subsets whose convex hulls intersect. Formally, given a set $P=\{x_{1},x_{2},\ldots,x_{d+2}\}\subseteq\mathbb{R}^{d}$, there exists a partition of $P$ into two disjoint subsets $P_{1}$ and $P_{2}$ ($P=P_{1}\cup P_{2}$) such that $\text{CH}(P_{1})\cap\text{CH}(P_{2})\neq\emptyset$, where here and in what follows $\text{CH}(X)$ denotes the convex hull of a set of points $X$. The bound $d+2$ is tight as one can easily see by taking any set of $d+1$ affine-independent points in $\mathbb{R}^{d}$, that is, the vertices of a $d$-dimensional simplex.
This fundamental theorem is a cornerstone in discrete geometry, providing insights into the structure of point sets and their convex combinations. It implies various fundamental theorems in geometry including the classical Helly’s Theorem, the center point theorem, and more. Its implications extend to machine learning, statistical learning and computational geometry, influencing algorithms for geometric separation, point location, and convex hull computations. An example of such implication is the existence of small so-called $\varepsilon$}-nets or $\varepsilon$}-approximations for range-spaces defined by semi-algebraic sets which are a core notion in those areas. The combinatorial notion of VC-dimension is, in fact, the analog of Radon’s bound for abstract set-systems.
An equivalent formulation of Radon’s theorem states that for any set $P$ of $d+2$ points in $\mathbb{R}^{d}$, there exists a partition of the set into two subsets $P_{1}$ and $P_{2}$ such that any convex set containing $P_{1}$ must intersect any convex set containing $P_{2}$. Radon’s theorem serves as a basis for numerous generalizations and related results, such as the following classical and beautiful theorem of Tverberg [19], which further enriches our understanding of geometric configurations.
Theorem 1.1.
(Tverberg’s Theorem[19]) Let $r\geq 2$ be a fixed integer and $d\geq 1$. Then for any set $P$ of (r1)(d+1)+1(r-1)(d+1)+1 points in $\mathbb{R}^{d}$ there exists a partition of $P$ into $r$ pairwise disjoint sets $P=\bigcup_{i=1}^{r}P_{i}$ such that $\bigcap_{i=1}^{r}\text{CH}(P_{i})\neq\emptyset$.
Note that Radon’s theorem is the special case of Tverberg’s theorem with $r=2$.
Tverberg’s Theorem has far-reaching implications in discrete and computational geometry and beyond. Combined with the colorful Carathéodory’s theorem of Bárány [5] it implies, for example, the so-called first selection Lemma (see, e.g., [14]).
An equivalent formulation of Tverberg’s theorem states that for any set $P$ of (r1)(d+1)+1(r-1)(d+1)+1 points in $\mathbb{R}^{d}$, there exists a partition of the set into $r$ pairwise disjoint sets $P=\bigcup_{i=1}^{r}P_{i}$ such that for any family of convex sets $C_{1},\ldots,C_{r}$ with $P_{i}\subset C_{i}$ for every $i\in[r]$ we have that $\bigcap_{i=1}^{r}C_{i}\neq\emptyset$.
In this paper we settle an open problem posed in the 1970s by Kalai and reiterated in later surveys (e.g., [6, 13]) that asks for a generalization of Radon’s Theorem (and Tverberg’s Theorem) to sets which are not necessarily convex nor even connected, but are the union of a bounded number of convex sets.
Definition 1.2.
Let $s\geq 1$ be an integer. A set $C$ in $\mathbb{R}^{d}$ is said to be $s$-convex if it is the union of $s$ convex sets.
Problem 1.3.
[[6]-problem 6.6] Determine or estimate the least integer $f=f(d,s,t)$ such that for any set $P$ of $f$ points in $\mathbb{R}^{d}$ there is a partition $P=A\cup B$ such that any $s$-convex set containing $A$ must intersect any $t$-convex set containing $B$.
Bárány and Kalai explain in [6] why $f(d,s,t)$ is finite using a Ramsey type argument. They mention that as is often the case when using hypergraph Ramsey Theory the upper bound obtained is horrible, and add that it would be nice to understand the correct behavior of this function. Despite substantial attention, this problem remained unsolved for more than fifty years.
A more general problem is the following:
Problem 1.4.
Determine or estimate the least integer $f=f_{r}(d,s_{1},\ldots,s_{r})$ such that for any set $P$ of $f$ points in $\mathbb{R}^{d}$ there is a partition into $r$ pairwise disjoint sets $P=\bigcup_{i=1}^{r}P_{i}$ such that for any family of sets $C_{1},\ldots,C_{r}$ with $P_{i}\subset C_{i}$ where $C_{i}$ is an $s_{i}$-convex set for every $i\in[r]$ we have that $\bigcap_{i=1}^{r}C_{i}\neq\emptyset$.
Notice that Radon’s theorem is equivalent to $f(d,1,1)=d+2$ and more generally Tverberg’s theorem is equivalent to $f_{r}(d,1,\ldots,1)=(r-1)(d+1)+1$.
1.1 The main results
Our first main result is a fairly simple proof of the following near-optimal upper bound:
Theorem 1.5.
$f(d,s,t)=O(dst\log(st+1))$
Our second main result is an extension of the theorem above to the following version of Tverberg’s theorem for unions of convex sets and any $r\geq 2$:
Theorem 1.6.
$f_{r}(d,s_{1},\ldots,s_{r})=O\left(dr^{2}\cdot\log r\cdot\Pi_{i=1}^{r}s_{i}\cdot{\ln(1+\Pi_{i=1}^{r}s_{i}})\right)$
Remark 1.7.
Following our paper, Chen et al. obtain in [9] nearly matching lower bounds for $d\geq 2r-2$. For this range of $d$ and $r\geq 2$ they prove that for all $s$,
$f_{r}(d,s,\ldots,s)\geq(d-2r+4)s^{r}.$
When each of the $s_{i}$-convex sets in the last theorem is a union of convex components with bounded overlap (specifically, $(\ell+1)$-wise disjoint), we obtain a sharper bound; see Theorem 5.3.
A notable feature of our proofs is that they are mainly combinatorial and therefore work for abstract separable convexity spaces with bounded Radon numbers, as we proceed to explain.
Definition 1.8.
An abstract convexity space is a pair $(X,\mathcal{C})$ where $X$ is a set and $\mathcal{C}\subseteq 2^{X}$ is a family of subsets of $X$ called convex sets, satisfying the following axioms:
  1. 1.
    $X\in\mathcal{C}$ and $\emptyset\in\mathcal{C}$.
  2. 2.
    $\mathcal{C}$ is closed under intersections, i.e., if $\mathcal{D}\subseteq\mathcal{C}$, then $\bigcap\mathcal{D}\in\mathcal{C}$.
This generalizes the notion of convex sets in $\mathbb{R}^{d}$. Abstract convexity spaces arise naturally in combinatorics, geometry, and lattice theory, and provide a unifying framework to study convexity. For more on abstract convexity spaces, see, e.g., [12, 15, 20].
For a subset $P\subset X$ in an abstract convexity space $(X,\mathcal{C})$ define its convex hull $\text{CH}(P)$ to be $\text{CH}(P)=\bigcap_{P\subset S,S\in\mathcal{C}}S$.
For a set $P\subset X$ we say that $P=P_{1}\cup P_{2}$ is a Radon partition if $\text{CH}(P_{1})\cap\text{CH}(P_{2})\neq\emptyset$. We say that an $r$ partition $P=\bigcup_{i=1}^{r}P_{i}$ is an $r$-Tverberg partition of $P$ if $\bigcap_{i=1}^{r}\text{CH}(P_{i})\neq\emptyset$. The Radon number $r(X,\mathcal{C})$ of a space $(X,\mathcal{C})$ is the minimum integer $n$ such that any subset $P\subset X$ with $\lvert{P}\rvert\geq n$ admits a Radon partition. If $n$ is unbounded then the Radon number is $\infty$. Similarly, the $r$’th Tverberg number $T_{r}(X,\mathcal{C})$ is the minimum integer $n$ such that any subset $P\subset X$ with $\lvert{P}\rvert\geq n$ admits an $r$-Tverberg partition.
A convex set $H\in\mathcal{C}$ in an abstract convexity space $(X,\mathcal{C})$ is called a halfspace if its complement is also a convex set i.e., $X\setminus H\in\mathcal{C}$.
We say that an abstract convexity space $(X,\mathcal{C})$ is separable if for any two disjoint sets there exists a separating halfspace. Formally, for any $C_{1}\in\mathcal{C}$ and $C_{2}\in\mathcal{C}$ if $C_{1}\cap C_{2}=\emptyset$ then there exists a halfspace $H$ such that $C_{1}\subset H,C_{2}\subset X\setminus H$.
Our proof technique of Theorem 1.5 and Theorem 1.6 works for abstract convexity spaces which are separable and have a bounded Radon number. Let $(X,\mathcal{C})$ be a separable abstract convexity space with Radon number $d=r(X,\mathcal{C})$. Call a subset $A\subset X$ $s$-convex if it is the union of $s$ convex sets in $\mathcal{C}$. Let $F=F(d,s,t)$ be the minimum integer so that any set $P$ of cardinality at least $F$ in a separable abstract convexity space $(X,\mathcal{C})$ with Radon number $d$ admits a partition $P=P_{1}\cup P_{2}$ such that any $s$-convex set containing $P_{1}$ must intersect any $t$-convex set containing $P_{2}$. We have:
Theorem 1.9.
$F(d,s,t)=O(dst\log{(st+1)})$.
More generally let $F=F_{r}(d,s_{1},\ldots,s_{r})$ be the minimum integer so that any set $P$ of cardinality at least $F$ in a separable abstract convexity space $(X,\mathcal{C})$ with Radon number $d$ admits a partition $P=\bigcup_{i=1}^{r}P_{i}$ into $r$ pairwise disjoint sets $P=\bigcup_{i=1}^{r}P_{i}$ such that for any family of sets $C_{1},\ldots,C_{r}$ with $P_{i}\subset C_{i}$ where $C_{i}$ is an $s_{i}$-convex set for every $i\in[r]$ we have that $\bigcap_{i=1}^{r}C_{i}\neq\emptyset$.
Theorem 1.10.
$F_{r}(d,s_{1},\ldots,s_{r})=O\left(dr^{2}\cdot\log r\cdot\Pi_{i=1}^{r}s_{i}\cdot{\ln(1+\Pi_{i=1}^{r}s_{i}})\right)$.
Theorem 1.10 for the special case $s_{i}=1$ for all $i$ provides an upper bound of $O(dr^{2}\log r)$ on the so-called Tverberg number of a separable abstract convexity space in terms of its Radon number $d$. This improves the upper bound of $c(d)r^{2}\log^{2}r$ by Bukh [8] who proved it for the more general setting of (not necessarily separable) abstract convexity spaces. In [17] Pálvölgyi provided an upper bound of the form $O(d^{d^{d^{\log d}}}r)$ which is linear in $r$ but super exponential in $d$.
We describe here the proofs of Theorem 1.5 and Theorem 1.6. Essentially the same proofs establish the corresponding results, Theorem 1.9 and Theorem 1.10, for abstract separable convexity spaces.
1.2 Structure
The rest of this paper is organized as follows. In Section 2 we provide background on VC-dimension and shatter functions, and introduce the $r$-VC-dimension. In Section 3 we describe the proof of Theorem 1.5. The proof of Theorem 1.6 is given in Section 4. After a discussion of lower and upper bounds for the relevant functions for some special cases in Section 5 we suggest a few open problems in the final Section 6.
2 Preliminaries: VC-dimension and $r$-VC-dimension
The Vapnik-Chervonenkis dimension of a hypergraph (which is defined below) is a measure of its complexity. This notion plays a central role in statistical learning, computational geometry, and other areas of computer science and combinatorics (see, e.g., [3, 2, 4, 7, 16]). Many graphs and hypergraphs that arise in geometry have bounded VC-dimension. Our proof of Theorem 1.5 is based on some simple properties of the VC-dimension and the shatter function of the relevant hypergraphs (also defined below). In order to prove Theorem 1.6 we define an extended version of the VC-dimension and show how to bound it in terms of the usual dimension.
Definition 2.1 (VC-dimension).
The Vapnik-Chervonenkis dimension $VC(H)$ of a hypergraph $H=(V,E)$ is the largest integer $d$ such that there exists a subset $S\subseteq V$ (not necessarily in $E$) with $|S|=d$ that is shattered by $E$. A subset $S$ is said to be shattered by $E$ if, for every subset $T\subseteq S$, there exists a hyperedge $e\in E$ such that $e\cap S=T$.
Remark 2.2.
Note that in a hypergraph $H=(V,E)$ a subset $S$ is shattered iff for every partition $S=A\cup B$ (with $A\cap B=\emptyset$) there exist two hyperedges $e_{1}\in E$ and $e_{2}\in E$ such that $A\subset e_{1}$, $B\subset e_{2}$ and $e_{1}\cap e_{2}\cap S=\emptyset$. We will need a generalization of this notion to partitions into $r>2$ sets. This generalization is given below in Definition 2.5.
[…]
Definition 2.3.
The primal shatter function of a hypergraph $H=(V,E)$ is the following function $\pi_{H}:\mathbb{N}\rightarrow\mathbb{N}$:
$\pi_{H}(m)=\max_{S\subseteq V,|S|=m}|\{S\cap e:e\in E\}|.$
The value $\pi_{H}(m)$ represents the maximum number of distinct subsets of a set $S$ of cardinality $m$ that can be realized as intersections with hyperedges in $E$.
The following lemma, known as the Sauer-Shelah-Perles lemma, provides an upper bound on the shatter function for hypergraphs with bounded VC-dimension (See, e.g., [14]):
Lemma 2.4 (Sauer-Shelah-Perles).
Let $H=(V,E)$ be a hypergraph with VC dimension $d$. Then
$\pi_{H}(m)\leq\sum_{i=0}^{d}\binom{m}{i}$.
In particular, if $m>d$, then $\pi_{H}(m)\leq(\frac{em}{d})^{d}$.
In order to tackle Problem 1.4 we need to develop an analogous notion of a shattered set in hypergraphs with bounded VC-dimension for partitions with more than $2$ parts. The relevant definition follows. The motivation for this definition will become clear from its application in the study of Problem 1.4
Definition 2.5 ($r$-VC-dimension).
Let $H=(V,E)$ be a fixed hypergraph. Fix an arbitrary set $S\subseteq V$ and consider a partition of $S$ into $r$ pairwise disjoint sets $S=S_{1}\cup\cdots\cup S_{r}$. We call such a partition realizable if there exist hyperedges $e_{1},\ldots,e_{r}\in E$ such that $S_{i}\subset e_{i}$ for all $i\in[r]$ and $S\cap\bigcap_{i=1}^{r}e_{i}=\emptyset$. A subset $S\subset V$ is said to be $r$-shattered by $E$ if any partition of $S$ into $r$ pairwise disjoint sets is realizable. For a Hypergraph $H=(V,E)$ and an integer $r\geq 2$ its $r$-VC-dimension is defined to be the supremum of the possible sizes of an $r$-shattered set in $H$.
Remark 2.6.
For any $2\leq r_{1}<r_{2}$ if a subset $S$ is $r_{1}$-shattered then it is also $r_{2}$ shattered. Indeed, consider any partition $S=A_{1}\cup A_{2}\cup\cdots\cup A_{r_{2}}$. Then by uniting $r_{2}-r_{1}+1$ sets of that partition we get an $r_{1}$ partition, say $S=A_{1}\cup\cdots\cup A_{r_{1}-1}\cup B$ for which there is a family of hyperedges $e_{1},\ldots,e_{r_{1}}$ ($A_{1}\subseteq e_{1},\ldots,A_{r_{1}-1}\subseteq e_{r_{1}-1},B\subseteq e_{r_{1}}$ for which $S\cap\bigcap_{i=1}^{r_{1}}e_{i}=\emptyset$ ) witnessing the partition which is realizable. The same family of hyperedges with $e_{r_{1}}$ taken $r_{2}-r_{1}+1$ times is a witness for the $r_{2}$ partition. Thus, the $r_{1}$-VC-dimension is at most the $r_{2}$-VC-dimension.
The following lemma provides an upper bound on the $r$-VC-dimension as a function of its classical VC-dimension:
Lemma 2.7.
There exists an absolute constant $C$ such that for every integer $d$ and any hypergraph $H=(V,E)$ with VC-dimension $d$ the following holds. For every integer $r\geq 2$, the $r$-VC-dimension is at most $Cdr^{2}\log r$. This bound is nearly optimal: for every $d$ and $r$ there is a hypergraph with VC-dimension $d$ that admits an $r$-shattered set of size $\Omega(dr^{2})$.
The proof of the upper bound in the lemma is described in Section 4. The proof of the lower bound is given implicitly in Theorem 5.6 as explained in the remark following it.
Even though we don’t use the following lemma in this paper we feel that it is worth mentioning as it provides an extension of the Sauer-Shelah-Perles Lemma for hypergraphs with $r$-VC-dimension $t$. The proof follows immediately from Corollary 1 in [1].
Lemma 2.8 ($r$-shatter-function).
Let $H=(V,E)$ be a hypergraph with $r$-VC-dimension $t$. For a subset $S\subset V$ let $\pi_{r}(S)$ denote the number of realizable $r$-partitions of $S$. Let $\pi_{r}(m)$ denote the maximum of $\{\pi_{r}(S)\mid|S|=m\}$. Then we have:
$\pi_{r}(m)\leq\sum_{i=0}^{t}{m\choose i}(r-1)^{m-i}\leq(r-1)^{m}\sum_{i=0}^{t}{m\choose i}$.
3 Proof of Theorem 1.5
Before proceeding to the proof of Theorem 1.5 we need the following two simple lemmas.
Lemma 3.1.
Let $C_{1}$ be an $s$-convex set in $\mathbb{R}^{d}$ and $C_{2}$ a $t$-convex set. Assume that $C_{1}\cap C_{2}=\emptyset$. Then there exist $s$ convex polytopes $K_{1},\ldots,K_{s}$ each having at most $t$ facets whose union covers $C_{1}$ so that the complement of the union covers $C_{2}$. Namely, $C_{1}\subset\bigcup_{i=1}^{s}K_{i}$ and $C_{2}\subset\overline{\bigcup_{i=1}^{s}K_{i}}$.
Proof.
Since $C_{1}$ is $s$-convex it can be written as $C_{1}=\bigcup_{i=1}^{s}X_{i}$ for some convex sets $X_{1},\ldots,X_{s}$. Similarly $C_{2}=\bigcup_{j=1}^{t}Y_{i}$ for $t$ convex sets $Y_{1},\ldots,Y_{t}$. Since $C_{1}\cap C_{2}=\emptyset$, for every $i\in[s],j\in[t]$ we have that $X_{i}\cap Y_{j}=\emptyset$ and therefore there exists a hyperplane $h_{i,j}$ strictly separating $X_{i}$ and $Y_{j}$. Assume without loss of generality that the positive open halfspace $h_{i,j}^{+}$ bounded by $h_{i,j}$ contains $X_{i}$ and the negative open halfspace $h_{i,j}^{-}$ contains $Y_{j}$. For every $i\in[s]$ let $K_{i}$ be the convex polytope which is the intersection $\bigcap_{j=1}^{t}h_{i,j}^{+}$. Note that $K_{i}$ is a convex polytope with at most $t$ facets containing $X_{i}$ and its complement $\overline{K_{i}}$ contains $C_{2}$. Thus the union of the polytopes $\bigcup_{i=1}^{s}K_{i}$ contains $C_{1}$ and its complement $\overline{\bigcup_{i=1}^{s}K_{i}}$ contains $C_{2}$. This completes the proof of the lemma. ∎
Lemma 3.2.
Let $l=st$ where $s,t\geq 1$ are integers, and let $H=(P,E)$ be a hypergraph where $P$ is a set of points in $\mathbb{R}^{d}$ and $S\in E$ is a hyperedge if and only if there exists a set $K_{1},\ldots,K_{s}$ of $s$ convex polytopes each having at most $t$ facets such that $S=P\cap(\bigcup_{j=1}^{i}K_{j})$. Equivalently, $S$ can be cutoff from $P$ by intersecting it with a set consisting of a union of $s$ convex polytopes each containing at most $t$ facets. Then the VC-dimension of $H$ is bounded by $O(dl\log(l+1))$.
Proof.
The proof is standard, hence we only sketch the argument. Suppose $X$ is a set of points. By Lemma 2.4 we have an upper bound for the number of intersections of $X$ with a halfspace. This number raised to the power $t$ is thus an upper bound for the number of distinct intersections of $X$ with a polytope having at most $t$ facets. Raising it again to the power $s$ we get an upper bound for the number of intersections of $X$ with a union of $s$ such polytopes. If $X$ is shattered then the last number should be at least $2^{|X|}$, providing the required bound. We omit the detailed computation. See also, e.g., [14](Proposition 10.3.3) for this reasoning. ∎
Proof of Theorem 1.5: Put $l=st$. Let $H=(\mathbb{R}^{d},E)$ be the hypergraph as in Lemma 3.2. Let $n=O(dl\log(l+1))=O(dst\log(st+1))$ be its VC-dimension. We claim that $f(d,s,t)\leq n+1$. Indeed, Let $P$ be a set of $n+1$ points in $\mathbb{R}^{d}$. Since $P$ cannot be shattered by the hyperedges in $H$ there exists a non-trivial subset $A\subset P$ such that no hyperedge $S\in E$ has the property that $S\cap P=A$. In other words, there does not exists a set $K$ which is the union of $s$ convex polytopes, each having at most $t$ facets such that $K\cap P=A$. We claim that the partition $P=A\cup(P\setminus A)$ has the property that every $s$-convex set containing $A$ must intersect any $t$-convex set containing $P\setminus A$. Indeed, assume to the contrary that there exists an $s$-convex set $C_{1}$ containing $A$ and a $t$-convex set $C_{2}$ containing $P\setminus A$ such that $C_{1}\cap C_{2}=\emptyset$. Then by Lemma 3.1 there exists a set $K$ which is the union of $s$ convex polytopes each having at most $t$ facets which contains $C_{1}$ with a complement $\overline{K}$ that contains $C_{2}$. In particular, $K$ contains $A$ and $\overline{K}$ contains $P\setminus A$ so $K\cap P=A$, a contradiction. This completes the proof. $\Box$
4 A Generalized Tverberg Theorem
In this section we tackle Problem 1.4. In what follows we provide a bound on $f_{r}(d,s_{1},\ldots,s_{r})$. To simplify the presentation we assume that $s_{1}=s_{2}=\cdots=s_{r}=s$ and abuse the notation writing $f_{r}(d,s)$ for $f_{r}(d,s,s,\ldots,s)$. Our proof technique can be easily modified to make the bound sensitive for any $r$ integer parameters $s_{1},\ldots,s_{r}$ in the more general setting.
Our argument is based on the notion defined in 2.5, which is an extension of the notion of VC-dimension to partitions with more than $2$ sets.
We need Lemma 2.7. We proceed with its proof.
Proof.
Let $r\geq 2$ be an integer. Suppose that
$\left(\sum_{i=0}^{d}{f\choose i}\right)^{r}<\left(\frac{r}{r-1}\right)^{f}$ (1)
We claim that then any subset $S$ of $f$ vertices cannot be $r$-shattered. Namely, there exists a partition $S=\bigcup_{i=1}^{r}S_{i}$ so that whenever we have $r$ hyperedges $e_{1},e_{2},\ldots,e_{r}\in E$ such that $S_{i}\subset e_{i}$ for all $i\in[r]$ it must hold that $S\cap\bigcap_{i=1}^{r}e_{i}\neq\emptyset$.
Note that the inequality (1) holds for $f=Cdr^{2}\log r$ for some absolute constant $C$ so any $r$-shattered set has size at most f1f-1.
To prove the claim suppose it is false and there is a set $S$ that violates the condition. For every (ordered) $r$-tuple of hyperedges $e_{1},e_{2},\ldots,e_{r}$ with no common intersection in $S$, every point of $S$ belongs to at most r1r-1 of these hyperedges. Therefore, these fixed $r$ hyperedges can be used to provide at most $(r-1)^{f}$ partitions into $r$ sets $S_{1},\ldots,S_{r}$. Indeed, each of the $f$ points of $S$ can be allocated to at most r1r-1 of the $r$ parts. If, for example, a point lies in the hyperedges $e_{1}$, $e_{2}$ and so on up to $e_{r-1}$, then in the partition this point can be in either $S_{1}$ or $S_{2}$ and so on up to $S_{r-1}$ but not in $S_{r}$, and the situation is similar in each other case. By the Sauer-Shelah-Perles Lemma 2.4, the number of ordered $r$-tuples of intersections of hyperedges with $S$ is at most
$\left(\sum_{i=0}^{d}{f\choose i}\right)^{r}$
Since we have to cover all $r^{f}$ ordered partitions of $S$ into $r$ parts we get
$\left(\sum_{i=0}^{d}{f\choose i}\right)^{r}\cdot(r-1)^{f}\geq r^{f}$
This contradicts the assumption (1) and completes the proof. ∎
We also need the following simple geometric result showing that $s$-convex sets with no common point can each be enclosed in an $s$-convex polytope (i.e., a union of $s$ convex polytopes), such that all those $s$-convex polytopes also have no common point. Moreover, we provide an upper bound on the number of facets of each of these polytopes.
Lemma 4.1.
Let $C_{1},C_{2},\ldots,C_{r}$ be $r$ sets in $\mathbb{R}^{d}$ where each set is an $s$-convex set. Assume that $\bigcap_{i=1}^{r}C_{i}=\emptyset$. Then there exist $r$ sets $K_{1},\ldots,K_{r}$ where each $K_{i}$ is the union of $s$ convex polytopes each having at most $s^{r-1}$ facets such that $C_{i}\subset K_{i}$ for all $i\in[r]$ and $\bigcap_{i=1}^{r}K_{i}=\emptyset$.
Proof.
Since $C_{i}$ is $s$-convex for any $i\in[r]$, it can be written as $C_{i}=\bigcup_{j=1}^{s}X_{i,j}$ for some convex sets $X_{i,1},\ldots,X_{i,s}$. Since $\bigcap_{i=1}^{r}C_{i}=\emptyset$, for every $i\in[r]$ we have that $C_{i}$ is disjoint from $B_{i}=\bigcap_{j\neq i}C_{j}$. Note that each such $B_{i}$ is the union of at most $s^{r-1}$ convex sets since it is an (r1)(r-1)-fold intersection of unions of $s$-convex sets. We construct the sets $K_{1},\ldots,K_{r}$ one by one. First, we replace $C_{1}$ by $K_{1}$ separating it from the union of at most $s^{r-1}$ convex sets $B_{1}=\bigcap_{j>1}C_{j}$. As before, this can be done with $K_{1}$ which is the union of $s$ convex polytopes, each having at most $s^{r-1}$ facets. In particular $K_{1}$ is also $s$-convex. Also $C_{1}\subset K_{1}$. Moreover, $K_{1}\cap\bigcap_{j>1}C_{j}=\emptyset$. We then apply the same argument to $C_{2}$ as the intersection of $K_{1},C_{2},..,C_{r}$ is empty and all the sets are $s$-convex. So we can find a set $K_{2}$ which is $s$-convex and consists of the union of $s$ convex polytopes each having at most $s^{r-1}$ facets and such that $C_{2}\subset K_{2}$ and $K_{1}\cap K_{2}\cap\bigcap_{j>2}C_{j}=\emptyset$. Continuing in the same manner we conclude that each $C_{i}$ can be replaced with such a $K_{i}$ so that $C_{i}\subset K_{i}$ for all $i\in[r]$, each $K_{i}$ is the union of $s$ convex polytopes, each with at most $s^{r-1}$ facets, and $\bigcap_{i=1}^{r}K_{i}=\emptyset$. This completes the proof of the lemma. ∎
A refinement of Lemma 4.1 under an additional bounded-overlap assumption on the convex components (namely, $(\ell+1)$-wise disjointness) is given in Section 5; see Theorem 5.3.
We are now ready to prove the following theorem extending Theorem 1.5 to partitions with $r>2$ parts, and generalizing Tverberg’s theorem to $s$-convex sets:
Proof of Theorem 1.6 (for $s_{1}=s_{2}=\ldots=s_{r}=s$).
Put $l=s^{r}$. Let $H=(\mathbb{R}^{d},E)$ be the hypergraph as in Lemma 3.2. Let $d^{\prime}=O(dl\log(l+1))$ be its VC-dimension. Let $n$ be the maximum size of an $r$-shattered set. Note that by Lemma 2.7 $n=O(d^{\prime}r^{2}\log r)=O(dr^{2}{\log r}\cdot s^{r}\cdot{\log(s^{r}+1)})$
We claim that $f_{r}(d,s)\leq n+1$. Indeed, let $P$ be a set of $n+1$ points in $\mathbb{R}^{d}$. Since $P$ cannot be $r$-shattered by the hyperedges in $H$ there exists a partition $P=\bigcup_{i=1}^{r}P_{i}$ such that whenever we have $r$ hyperedges $e_{1},\ldots,e_{r}\in E$ with $P_{i}\subset e_{i}$ for each $i\in[r]$ it follows that $P\cap\bigcap_{i=1}^{r}e_{i}\neq\emptyset$. In other words there do not exist sets $K_{1},\ldots,K_{r}$ each of which is the union of $s$ convex polytopes, each having at most $s^{r-1}$ facets, such that $P_{i}\subset K_{i}$ for every $i\in[r]$ and $\bigcap_{i=1}^{r}K_{i}=\emptyset$. We claim that the partition $P=\bigcup_{i=1}^{r}P_{i}$ has the property that for every family of $r$ sets $C_{1},\ldots,C_{r}$ such that for each $i\in[r]$ $P_{i}\subset C_{i}$ and each $C_{i}$ is an $s$-convex set it must hold that $\bigcap_{i=1}^{r}C_{i}\neq\emptyset$. Indeed, assume to the contrary that there exist $C_{1},\ldots,C_{r}$ such that for each $i\in[r]$ $P_{i}\subset C_{i}$, each $C_{i}$ is an $s$-convex set and $\bigcap_{i=1}^{r}C_{i}=\emptyset$. Then by Lemma 4.1 there exist sets $K_{1},\ldots,K_{r}$ for which $C_{i}\subset K_{i},\forall i\in[r]$ and each $K_{i}$ is the union of $s$ convex polytopes each having at most $s^{r-1}$ facets and $\bigcap_{i=1}^{r}K_{i}=\emptyset$, a contradiction. This completes the proof. ∎
Remark 4.2.
Chen et al. [9] showed that in the special case where the $s$-convex sets in the theorem are required to be pairwise disjoint then one can improve the bound in Theorem 1.6 to $C_{d,r}s^{2d+3}$ where $C_{d,r}$ is a constant depending only on $d$ and $r$. We can further improve this bound to $C_{d,r}s^{d+1}\log s$ and generalize it to the case where the $s$-convex sets are not necessarily pairwise disjoint but only $l+1$-wise disjoint, meaning that any $l+1$ sets of them are disjoint for some fixed $1\leq l$. The proof is essentially the same, replacing the bound for the number of facets of each polytope in $K_{i}$ by $O_{l,d,r}(s^{d})$ as described in Theorem 5.3.
5 Improved bounds in special cases
As before we denote the function $f_{r}(d,s_{1},s_{2},\ldots,s_{r})$ in which $s_{i}=s$ for all $i$ by $f_{r}(d,s)$.
In this section we discuss improved upper and lower bounds for the functions $f$ and $f_{r}$ in several special cases.
5.1 The Asymmetric Case $f(d,s,1)$
We start with the following asymmetric case providing the exact value of $f(2,s,1)$, sharp asymptotic bounds on $f(d,s,1)$ for $d\geq 4$ and near optimal upper bound on $f(3,s,1)$:
Theorem 5.1.
$f(2,s,1)=2s+2$
and for every $d\geq 4$
$f(d,s,1)=\Theta(ds\log(s+1))$
Proof.
Let $l$ denote the VC-dimension of a hypergraph defined by points in $\mathbb{R}^{d}$ with respect to convex polytopes with at most $s$ facets. We first show that $f(d,s,1)=l+1$. Combined with the known bounds on the VC-dimension of such hypergraphs we get the claimed bounds. If $l$ is the VC-dimension of such a hypergraph then there exists a shattered set $P$ of size $l$. Then for any partition $P=P_{1}\cup P_{2}$ there is a polytope $K$ with at most $s$ facets such that $P_{2}\subset K$ and $K\cap P_{1}=\emptyset$. In particular $P_{1}$ is contained in the union of the at most $s$ complement halfspaces supporting the facets of $K$, which is an $s$-convex set (consisting of the at most $s$ complement halfspaces), and $P_{2}$ lies in its complement. Since this holds for any partition it follows that $f(d,s,1)\geq l+1$. In order to show equality we need to show that every $l+1$ point set in $\mathbb{R}^{d}$ admits a partition that cannot be realized by intersections with disjoint convex and $s$-convex sets. This follows by the same argument from the fact that any set $P$ of $l+1$ points cannot be shattered in the corresponding hypergraph and therefore there is at least one partition $P=P_{1}\cup P_{2}$ so that no polytope with at most $s$ facets can contain $P_{2}$ while being disjoint from $P_{1}$. Hence any $s$-convex set $C$ containing $P_{1}$ must intersect the convex hull $\text{CH}(P_{2})$ for otherwise by separation arguments as above there would be a polytope which is the intersection of $s$ half spaces containing $P_{2}$ and disjoint from $P_{1}$, a contradiction. It is a simple exercise to see that in the plane ($d=2$), the VC-dimension is $l=2s+1$ and thus $f(2,s,1)=l+1=2s+2$. The statement for any fixed $d\geq 4$ follows from the results in [10]. ∎
Theorem 5.2.
$f(3,s,1)\leq 4s+1$
Proof.
We need to prove that for every set $P$ of $4s+1$ points there exists a partition $P=A\cup B$ such that any convex set containing $A$ must intersect any $s$-convex set containing $B$. Notice that by the above arguments it is enough to prove that the VC-dimension of the hypergraph $H=(P,E)$ where $E$ is the family of all intersections of $P$ with a convex polytope with at most $s$ facets is bounded by $4s$. This fact is proved in [11]. For completeness we include a proof based on an argument in [18]. Assume to the contrary that there exists a set $P\subset\mathbb{R}^{3}$ of size $4s+1$ that is shattered by $H$. In particular, for any partition $P=A\cup B$ if $A$ can be separated from $B$ by a convex polytope with at most $s$ facets, then there exists a set of $s$ half spaces whose union contains $B$ but none of the points in $A$.
Next, we need the following fact which was proved in [18]: There exists a $4$-coloring of the points of $P$ such that no halfspace that contains at least two points of $P$ is monochromatic.
Consider such a coloring. By the pigeonhole principle there is a monochromatic set $B\subset P$ of size at least $s+1$. We claim that $B$ cannot be separated from $A=P\setminus B$ with a union of only $s$ halfspaces. Indeed, if such $s$ halfspaces exist then one of them must contain at least $2$ points of $B$ and none of the points in $A$ so such a halfspace cuts off a monochromatic set of points, a contradiction. This completes the proof.
5.2 Upper Bounds
In the special case where each of the sets $C_{i}$ in Lemma 4.1 above has the property that the $s$ convex sets $X_{i,1},\ldots,X_{i,s}$ whose union is $C_{i}$ are in addition $l+1$-wise disjoint (any intersection of $l+1$ of them is empty) then one can make sure that each polytope defining the sets $K_{i}$ has at most $O_{r,d,l}(s^{d})$ facets.
Theorem 5.3.
Let $r\geq d+1$ be an integer and $C_{1},C_{2},\ldots,C_{r}$ be $r$ sets in $\mathbb{R}^{d}$ where each set is an $s$-convex set. Assume that $\bigcap_{i=1}^{r}C_{i}=\emptyset$ and that each $C_{i}$ is the union of s convex sets which are $l+1$-wise disjoint. Then there exist $r$ sets $K_{1},\ldots,K_{r}$ where each $K_{i}$ is the union of $s$ convex polytopes each having at most $O_{r,d,l}(s^{d})$-facets such that $C_{i}\subset K_{i}$ for all $i\in[r]$ and $\bigcap_{i=1}^{r}K_{i}=\emptyset$.
Proof.
We follow the notations as in the proof of Lemma 4.1. The main idea is to use the following known fact: If r1r-1 convex sets $X_{1,j_{1}},\ldots,X_{r-1,j_{r-1}}$ in $\mathbb{R}^{d}$ (for $r-1\geq d$) have a non-empty intersection then there is a subset of at most $d$ indices $I\subset[r-1]$ such that the lexicographic minimum point of $\cap_{i\in I}X_{i,j_{i}}$ equals the lexicographic minimum point of $\cap_{i=1}^{r-1}X_{i,j_{i}}$ (see e.g., Lemma 8.1.2 in [14]). Thus, consider for example the $s$-convex set $K_{r}$ constructed in Lemma 4.1 which is a union of $s$ convex polytopes each being the intersection of at most $s^{r-1}$ halfspaces. We can reduce the number of halfspaces as follows: The intersection $\cap_{i=1}^{r-1}C_{i}$ consists of a family of intersections of r1r-1 convex sets $X_{1,j_{1}},\ldots,X_{r-1,j_{r-1}}$. Each such nonempty intersection can be charged to some lexicographic minimum point of the intersection of a subfamily $\{X_{i,j_{i}}|i\in I\}$ for some $I\subset[r-1]$ with $|I|=d$. For any such fixed set $I\subset[r-1]$ with $|I|\leq d$, there are at most $l^{(r-1-|I|)}$ (r1r-1)-tuples of convex sets that charge such a point since this point can belong to at most $l$ convex sets of the form $X_{i,j_{i}}$ for any $i\in[r-1]\setminus I$. Clearly the intersection of each such tuple is convex and there are at most $\sum_{i=0}^{d}{r-1\choose i}s^{i}l^{r-1-i}=O_{r,d,l}(s^{d})$ such tuples. Thus, for each of the convex components $X_{r,i}$ $i\in[s]$ of $C_{r}$ we can use at most these many halfspaces. Altogether $K_{r}$ consists of $s$-convex polytopes, each having at most $O_{r,d,l}(s^{d})$ facets. ∎
The following simple result shows that the lower bound for $f_{r}(d,s)$ proved in Theorem 5.6 is tight up to a constant factor in dimension $d=1$.
Theorem 5.4.
$f_{r}(1,s)\leq r(r-1)(s+1)+1$
Proof.
Put n=r(r1)(s+1)+1n=r(r-1)(s+1)+1 and let $0<p_{1}<p_{2}<\ldots<p_{n}<1$ be a set of $n$ points on the line. We have to show that there is a coloring of these points by $r$ colors $0,1,2,\ldots,r-1$, so that for any $r$ sets $C_{i}$, where each $C_{i}$ is a union of at most $s$ intervals that covers all points of color $i$, there is a point that lies in all sets $C_{i}$. Naturally, the coloring we choose colors the points periodically, that is, the color of $p_{i}$ is defined to be $i\bmod r$. Let $C_{i}$ be collections of intervals as above. Note that each $C_{i}$ must contain all points of $P$ besides at most (s+1)(r1)(s+1)(r-1). Indeed, since it contains all points colored $i$, the gap between any two consecutive intervals in it contains at most r1r-1 points, and the same holds for the gap between 0 and its leftmost point and between $1$ and its rightmost point. (We note that here we can slightly improve the bound since, for example, the gap between 0 and the leftmost point of $C_{i}$ can contain at most $i$ points $p_{i}$). It follows that if n>r(r1)(s+1)n>r(r-1)(s+1) then there is a point (of $P$, although that’s not needed) that belongs to all sets $C_{i}$. This completes the proof. ∎
5.3 Lower Bounds
Remark 5.5.
A trivial lower bound for the function $f_{r}(d,s)$ can be proved by taking $s$ translated copies of an extremal example for the classical theorem of Tverberg with pairwise disjoint convex hulls. This gives the following.
$f_{r}(d,s)>s(d+1)(r-1).$
This is, of course, tight for $s=1$, as Tverberg’s Theorem is tight.
The next result shows that for $s\geq 3$ and any dimension $d\geq 1$, the lower bound becomes quadratic in $r$. Note that for $d\geq 2r-2$ a stronger bound of $(d-2r+4)s^{r}$ is proved in [9], but the bound below holds for all $d$. As shown in Theorem 5.4 above, for $d=1$ the lower bound is tight up to a constant factor for all $r$ and $s\geq 3$. For convenience we describe the proof for even $r$, a similar bound for odd $r$ follows from the one for r1r-1.
Proposition 5.6.
For every $d\geq 1,s\geq 3$ and even $r\geq 2$,
$f_{r}(d,s)>\frac{1}{4}\left(\left\lfloor\frac{d}{2}\right\rfloor+1\right)\left\lfloor\frac{s-1}{2}\right\rfloor r^{2}$
Proof.
Put $m=(\lfloor\frac{d}{2}\rfloor+1)r/2$, $p=\lfloor\frac{s-1}{2}\rfloor r/2$ and $n=mp$. Let $P$ be a set of $n$ points on the moment curve in $\mathbb{R}^{d}$ consisting of the $n$ points $z_{i}=(t_{i},t_{i}^{2},\ldots t_{i}^{d})$, where $0<t_{1}<t_{2}<\ldots<t_{n}<1$. Let $I_{1},I_{2},\ldots,I_{p}$ be $p$ open intervals that cover $(0,1)$ and appear on it in this order, where the left endpoint of each interval $I_{j+1}$ is just slightly smaller than the right endpoint of $I_{j}$ for each $j$. The intervals are chosen so that each $t_{i}$ belongs to exactly one such interval, and each interval contains exactly $m$ points $t_{j}$.
In order to complete the proof we show that for any coloring of the points of $P$ by $r$ colors there are $s$-convex sets $C_{i}$, with $C_{i}$ containing all points of color $i$, so that the intersection of all the sets $C_{i}$ is empty. Each set $C_{i}$ will be defined as the union of at most $s$ convex sets, where each of the convex sets will be the convex hull of points of color $i$ with first coordinate lying in a union of some consecutive intervals $I_{q}$. The crucial property of the definition of these convex sets is that for each interval $I_{q}$ there will be an index $i$ so that one of the convex sets in the definition of $C_{i}$ will be the set of all points $z_{j}=(t_{j},t_{j}^{2},\ldots,t_{j}^{d})$ of color $i$ for which $t_{j}$ lies in this single interval $I_{q}$. Moreover, this will be done in a way that ensures that the number of such points is always at most $\lfloor d/2\rfloor$.
Note that if such a choice indeed exists, then the intersection of the corresponding $r$ sets $C_{i}$ will be empty as needed. Indeed, otherwise the intersection is some point $z\in\mathbb{R}^{d}$ (not necessarily on the moment curve). Let $t$ denote the first coordinate of this point $z$. Suppose first that $t$ belongs only to one interval $I_{q}$, and let $i$ be the color chosen for $I_{q}$ so that there are at most $\lfloor d/2\rfloor$ points of color $i$ with their first coordinate in $I_{q}$. Then $z$ has to lie in the convex hull of these points (as any other convex set among the ones defining $C_{i}$ either has all its points with first coordinate smaller than $t$ or all its points with first coordinate larger than $t$). Since any finite subset of the moment curve is $\lfloor d/2\rfloor$-neighborly [^1] A set of points is $r$-neighborly if any subset of at most $r$ points form a face of the convex-hull of the set., this convex hull is disjoint from the convex hull of all the other points of $P$, implying that $z$ cannot lie in any other set $C_{g}$ besides $C_{i}$. If $t$ belongs to two intervals, say $I_{q-1}$ and $I_{q}$ and $i$ is defined as before, then $z$ cannot lie in $C_{i}$, since in this case any convex set among those defining $C_{i}$ either has all its points with first coordinate smaller than $t$ or all its points with first coordinate larger than $t$.
It thus only remains to show that we can choose for every interval $I_{q}$ a color $i$ with the required properties, making sure that no fixed color is chosen more than $\lfloor(s-1)/2\rfloor$ times (as this way each set $C_{i}$ will be the union of at most $s$ convex sets). To do so we go over the intervals $I_{q}$ one by one, in an arbitrary order (for example, from left to right). When dealing with the interval $I_{q}$ after handling several previous ones, we first note that since there are exactly $m=(\lfloor d/2\rfloor+1)r/2$ points with first coordinate in $I_{q}$, there are at least $r/2$ colors that appears at most $\lfloor d/2\rfloor$ times in $I_{q}$. As so far we have selected the colors $i$ for at most $p-1=\lfloor(s-1)/2\rfloor r/2-1$ intervals, there are less than $r/2$ colors $i$ that have already been chosen $\lfloor(s-1)/2\rfloor$ times. It follows that there is at least one color $i$ that can be chosen for the interval $I_{q}$, implying that the process terminates successfully. This completes the proof of the theorem. ∎
Remark 5.7.
Note that for $d=1$, any hypergraph whose vertices are an arbitrary set of points on the line, and whose hyperedges are the intersections of this set of points with union of $s$ intervals, has VC-dimension at most $2s$. Indeed, it is easily verified that no set of size $2s+1$ is shattered. Assume to the contrary that there is a set $P=\{x_{1},\ldots,x_{2s+1}\}$ that is shattered where the points are indexed by their increasing order along the line. Notice that every union of $s$-intervals that contains only the $s+1$ points with odd indices $\{x_{1},x_{3},\ldots,x_{2s+1}\}$ must have one of its intervals containing two consecutive such points and hence also a point with an even index, a contradiction. Therefore, Theorem 5.6 in dimension $1$ provides a hypergraph with VC-dimension bounded by $D=2s$ and an $r$-shattered set of size at least
$\frac{1}{4}\left\lfloor\frac{s-1}{2}\right\rfloor r^{2}=\frac{1}{4}\left\lfloor\frac{D-2}{4}\right\rfloor r^{2}.$
This establishes the claimed lower bound in Lemma 2.7
6 Concluding remarks and open problems
We established extensions of Radon’s Theorem and Tverberg’s Theorem for unions of convex sets. The main tools in the proofs are upper bounds for the shatter functions of range spaces with a bounded VC-dimension as well as an extension of these results to $r$-VC-dimension. This extension, defined and studied here, is useful in the study of partitions with more than $2$ parts, which are the ones considered in the classical definition of the VC-dimension.
Our notion of $r$-VC-dimension may be of independent interest beyond the geometric problems studied here. It controls the growth of realizable $r$-partitions (a natural multiclass analogue of the usual shatter function), and thus provides a convenient complexity measure when one needs uniform bounds simultaneously over many parts. We expect that this parameter may be useful in other settings like the study of realizability of $r$-labelings or $r$-way classification rules in machine learning.
As mentioned in the introduction, already for $s_{i}=1$ for all $i$ the upper bound provided in our Theorem 1.10 is $O(dr^{2}\log(r+1))$. It improves the upper bound of $c(d)r^{2}\log^{2}(r+1)$ by Bukh [8] who proved it for the more general setting of (not necessarily separable) abstract convexity spaces. In [17] Pálvölgyi provided an upper bound of the form $O(d^{d^{d^{\log d}}}r)$ which is linear in $r$ but super exponential in $d$. It will be interesting to decide if one can get rid of the separability assumption and prove the bound in Theorem 1.10 for all abstract convexity spaces.
While our upper bounds and the lower bounds in [9] for the functions $f(d,s,t)$ and $f_{r}(d,s_{1},s_{2},\ldots,s_{r})$ are not very far from each other, the problem of determining them precisely remains open for most values of the parameters. It will be interesting to close the gap in these bounds.
Acknowledgements
We thank Gil Kalai for fruitful discussions.